# Pychop: Emulating Low-Precision Arithmetic in Numerical Methods and Neural Networks

### v4

@@ 3. The Pychop library
Numbers in machines are often approximated with truncation using discrete representations due to the constraints arisen from finite storage. Two fundamental representations dominate this domain: floating-point representation, which excels in representing a wide range of values with variable precision, and fixed-point representation, which prioritizes simplicity and efficiency in constrained environments. This section formulates both representations, with a particular focus on the IEEE 754 standard for floating-point representation, and compares their theoretical and practical implications in computational tasks. Our mixed-precision emulation software is implemented as a modular Python code supported by the backends of NumPy, PyTorch, and JAX, and it comprises three primary components: the Chop, Chopf, and Chopi class for float point, fixed point, and integer quantization, respectively, which are also integrated in the layers class with straight-through estimator for neural network deployment, supporting emulation for various representation format (see Table 3 for details) as well as subnormal numbers. In the following, we detail its design and implementation, emphasizing its main principles and user-friendly features.
Table 3. Rounding Modes for the rmode in low precision emulation.
@@
3 Round toward $\infty$- ∞ (round down) Rounds to the largest representable value less than or equal to the input, directing towards negative infinity.
@@ 3.1. Floating-point arithmetic
The IEEE 754 standard defines several formats, with the most common being single precision (32 bits) and double precision (64 bits). For a format with a $k$-bit exponent and a $p-1$-bit mantissa fraction, the bit layout is
$[s\,|\,e\,|\,m],$
where $s$ is the 1-bit sign, $e$ is the $k$-bit exponent field, and $m$ is the $p-1$-bit fractional part of the mantissa—the total precision $p$ includes an implicit leading bit for normalized numbers.
@@
  • Normalized Numbers ($21\leq e\leq 2^{k}-$ 2): The mantissa is $M=1+\sum_{i=1}^{p-1}m_{i}\cdot 2^{-i}$, where $m_{i}$ are the bits of the fractional field, and the exponent is $E=e-\text{bias}$, with $\text{bias}=2^{k-1}-1$. Thus
@@ 3.2. Implementation details
To round to low precision binary floating-point format, Pychop provides two modules, namely Chop and LightChop. Chop features a greater set of functionalities and follows the design of MATLAB chop; LightChop is deigned to be a lightweight “Chop” with essential features and enables efficient vectorization. Different backends of LightChop offer different features. The Torch and JAX backend offers GPU deployment, while the NumPy backend leverages Dask (Rocklin, 2015), a parallel computing library, to process a large NumPy array by chunking the array with user-defined chunk size, where the chunk size dictates how the array is split into smaller, manageable chunks for parallel or out-of-core computation, enabling efficient, scalable processing through lazy evaluation and a task graph that executes only when triggered, balancing memory use and parallelism based on the chosen chunk size.
The calling of Pychop is straightforward. In the following, we will illustrate how to use Chop and LightChop to emulate low precision arithmetic. The prototype of Chop and LightChop are separately described as below:
@@
  • 1LightChop(exp_bits:=int, sig_bits:=int, chunk_size:=int=1000, rmode:=int=1, subnormal:=bool=True, chunk_size: =int=1000, random_state:=int=42)
Generally, LightChop is faster than Chop, which will be verified in Section 4, but Chop includes more floating-point emulation features. We demonstrate their basic usage below:
1from Pychop import Chop
2import numpy as np
3
4X = np.random.randn(5000, 5000)
5ch = Chop(’h’, rmode=1) # Standard IEEE 754 half precision
6# other parameters are left as default.
7Xq = ch(X) # Rounding values
Use built-in precision
1import numpy as np
2from Pychop import Chop, LightChop
3from Pychop import Customs
4
5X = np.random.randn(5000, 5000)
6ch = Chop(customs=Customs(exp_bits=5, sig_bits=10), rmode=1) # half precision (5 exponent bits, 10+(1) significand bits, (1) is implicit bits)
7Xq = ch(X)
8
9ch = LightChop(exp_bits=5, sig_bits=10, rmode=1)
10Xq = ch(X)
Customized precision
@@ 3.3. Fixed-point arithmetic
Fixed-point is associated with numbers of a fixed number of bits, allocating a predetermined number of bits to the integer part and the fractional part, unlike floating-point representations that use an exponent and significand. The shift to fixed-point arithmetic is driven by several key advantages. First, fixed-point compute units are typically faster and significantly more efficient in terms of hardware resources and power consumption compared to floating-point units. Their smaller logic footprint allows for a higher density of compute units within a given area and power budget. Second, low-precision data representation reduces the memory footprint, enabling larger models to fit within available memory while also lowering bandwidth requirements. Fixed-point operations align well with digital signal processors (DSPs) and field-programmable gate arrays (FPGAs) because many lack dedicated floating-point units or optimize for fixed-point arithmetic. Collectively, these benefits enhance data-level parallelism, leading to substantial performance gains (Gupta et al., 2015).
@@
where $I$ is an integer, which may be signed (typically in two’s complement) or unsigned, $f$ is the fixed number of fractional bits, and $n$ is the total number of bits is, with $n-f$ bits allocated to the integer part and $f$ bits to the fractional part.
@@ 3.3.1. Process of fixed-point quantization
where the ibits refers to the bitwidth of integer part, fbits refers to the bitwidth of fractional part, and rmode indicates the rounding mode used in (3.3.1); the supported rounding modes can be found in Table 3.
The usage is given by the following example.
1from Pychop import Chopf
2import numpy as np
3
4X = np.random.randn(5000, 5000)
5ch = Chopf(ibits=4, fbits=4, rmode=1)
6ch(X)
Fix-point representation
3.3.2. Relation to and Deep Learning Deployment
@@ 3.4. Integer arithmetic
Integer quantization is a fundamental technique in digital systems to approximate real numbers using a finite set of integers, enabling efficient storage and computation in applications such as machine learning, digital signal processing, and embedded systems. Two primary approaches to quantization exist: symmetric quantization, which balances positive and negative ranges around zero, and asymmetric quantization, which allows unequal ranges for positive and negative values. In principle, integer quantization maps a real number $r\in\mathbb{R}$ to a discrete integer value $x\in\mathbb{Z}$ through scaling and rounding. The process can be adapted for either symmetric or asymmetric quantization depending on the range and scaling strategy. Consider a real number $r$ within a specified range $[r_{\text{min}},r_{\text{max}}]$. The goal is to represent $r$ using $n$-bit integers, defining a discrete set of quantization levels. The general quantization process involves scaling, rounding, and clamping, with variations depending on whether symmetric or asymmetric quantization is used.
@@
    For an $n$-bit signed integer in two’s complement, the representable range is $[-2^{n-1},2^{n-1}-1]$. The quantization range is defined symmetrically as $[-\omega,\omega]$, where $\omega$ is the maximum absolute value to be represented. The scaling factor $\Delta$ is:
@@
    $\overline{x}=\mathrm{R}(s),\quad\text{ where }\mathrm{R}(\cdot)\text{ denotes general rounding operator.}$.
@@
  • Asymmetric Quantization: Asymmetric quantization allows unequal ranges for positive and negative values, typically by defining a range $[r_{\text{min}},r_{\text{max}}]$ where $r_{\text{min}}\neq-r_{\text{max}}$. This is useful when the data distribution is skewed (e.g., all positive values in rectified activations like ReLU).
    For an $n$-bit signed integer, the range is still $[-2^{n-1},2^{n-1}-1]$, but the quantization maps $[r_{\text{min}},r_{\text{max}}]$ to this range. The scaling factor $\Delta$ is
@@
    $\overline{x}$ $=\mathrm{R}(s),\quad\text{ where }\mathrm{R}(\cdot)\text{ denotesgeneral rounding operator,}$,
@@
    The Chopi framework offers tailored integer quantization, allowing users to specify the bitwidth length (bits) and choose between symmetric or asymmetric quantization (symmetric). Additionally, two channel-specific parameters enable further customization: per_channel determines whether quantization is applied on a per-channel basis, while axis specifies the dimension along which channel-wise quantization occurs when is set to True. Below is a simple demonstration:
    1from Pychop import Chopi
    2
    4X_q = ch.quantize(X_np) # to integers
    5X_inv = ch.dequantize(X_q) # back to floating-point numbers
@@ 3.5. Common mathematical functions support and array manipulation routines
We simulate common mathematical functions and operations (such as built-in functions in NumPy, PyTorch, or JAX) in low-precision arithmetic by first rounding the input to low precision, performing operations in the working precision, and then rounding the result back to low precision. This approach contrasts with CPFloat, which applies mathematical operations in working precision to inputs in working precision before rounding the final result to low precision.
Table 4. Functions Support and Array Manipulation Routines (Part 1)
33footnotemark: 3
@@
sin Computes sine. Input in radians; output in $[-1,1]$.
cos Computes cosine. Input in radians; output in $[-1,1]$.
@@
arcsin Computes arcsin. Input in $[-1,1]$; output in $[-\pi/2,\pi/2]$ radians.
arccos Computes arccos. Input in $[-1,1]$; output in $[0,\pi]$ radians.
arctan Computes arctan. Output in $[-\pi/2,\pi/2]$ radians.
@@
tanh Computes hyperbolic tangent. Output in $(-1,1)$.
@@
arccosh Computes inverse hyperbolic cosine. Input $\geq 1$; output in $[0,\infty)$.
arctanh Computes inverse hyperbolic tangent. Input in $(-1,1)$; output real.
@@
log1p Computes $\log(1+x)$. Input $>-1$; enhanced precision for small $x$.
@@
erf Computes error function. Output in $(-1,1)$.
@@
Table 5. Functions Support and Array Manipulation Routines (Part 2)
 [^4] All functions are computed with chopping to enforce low-precision format, where applicable.
@@
sign Computes sign. Returns $-1$, 00, or $1$.
clip Clips values to range $[a_{\text{min}},a_{\text{max}}]$.
@@
The usage of common functions is illustrate as an example below:
1ch = Chop(’q43’, rmode=1)
2ch.sin(X)

### v5

@@ 3. The Pychop library
Numbers in machines are often approximated with truncation using discrete representations due to the constraints arisen from finite storage. Two fundamental representations dominate this domain: floating-point representation, which excels in representing a wide range of values with variable precision, and fixed-point representation, which prioritizes simplicity and efficiency in constrained environments. This section formulates both representations, with a particular focus on the IEEE 754 standard for floating-point representation, and compares their theoretical and practical implications in computational tasks. Our mixed-precision emulation software is implemented as a modular Python code supported by the backends of NumPy, PyTorch, and JAX, and it comprises three primary components: the FaultChop, Chopf, and Chopi class for float point, fixed point, and integer quantization, respectively, which are also integrated in the layers class with straight-through estimator for neural network deployment, supporting emulation for various representation format (see Table 3 for details) as well as subnormal numbers. In the following, we detail its design and implementation, emphasizing its main principles and user-friendly features.
Table 3. Rounding modes for the rmode in reduced-precision emulation.
@@
3 Round toward -$\infty$ (round down) Rounds to the largest representable value less than or equal to the input, directing towards negative infinity.
@@ 3.1. Floating-point Arithmetic
The IEEE 754 standard defines several formats, with the most common being single precision (32 bits) and double precision (64 bits). For a format with a $k$-bit exponent and a p1p-1-bit mantissa fraction, the bit layout is
[s|e|m],[s\,|\,e\,|\,m],
where $s$ is the 1-bit sign, $e$ is the $k$-bit exponent field, and $m$ is the p1p-1-bit fractional part of the mantissa—the total precision $p$ includes an implicit leading bit for normalized numbers.
@@
  • Normalized Numbers ($1\leq e\leq 2^{k}-2$): The mantissa is $M=1+\sum_{i=1}^{p-1}m_{i}\cdot 2^{-i}$, where $m_{i}$ are the bits of the fractional field, and the exponent is $E=e-\text{bias}$, with $\text{bias}=2^{k-1}-1$. Thus
@@
In the following, we present two fundamental modules that Pychop offers—namely Chop and FaultChop—for rounding numbers to reduced-precision binary floating-point format. FaultChop features a greater set of functionalities corresponding to Nick Higham’s original implementation [^3] https://github.com/higham/chop, with minor code optimizations for improved performance. In contrast to Chop, FaultChop retains full support for soft error simulation, whereas Chop simplifies certain operations to achieve greater speedup. In other words, Chop is designed to be a lightweight “FaultChop” with essential features and enables efficient vectorization. Generally, Chop is faster than FaultChop, which will be verified in Section 4, but FaultChop includes more floating-point emulation features. As such, Chop is designed as a high-performance interface optimized for neural network training and large-scale arrays, while FaultChop is designed as a full-featured research interface supporting complete custom formats and soft-error simulation via random bit-flipping in the significand (flip=True). This makes it particularly suitable for studying numerical robustness under hardware faults. We demonstrate their basic usage in Appendix A.1.
Further, different backends of Chop offer different features. The Torch and JAX backends offer GPU deployment, while the NumPy backend leverages Dask (Rocklin, 2015), a parallel computing library, to process a large NumPy array by chunking the array with user-defined chunk size, where the chunk size dictates how the array is split into smaller, manageable chunks for parallel or out-of-core computation, enabling efficient, scalable processing through lazy evaluation and a task graph that executes only when triggered, balancing memory use and parallelism based on the chosen chunk size.
The calling of Pychop is straightforward. In the following, we will illustrate how to use Chop and FaultChop to emulate reduced-precision arithmetic. The prototype of Chop and FaultChop are separately described as below:
  • 1Chop(exp_bits:=int, sig_bits:=int, chunk_size:=int=1000, rmode:=int=1, subnormal:=bool=True, chunk_size: =int=1000, random_state:=int=42)
@@
We intentionally provide full support for subnormal numbers (with a user-controllable switch to enable or disable them) as a core capability, primarily to allow Pychop to faithfully emulate the complete semantics of the IEEE 754 standard, rather than merely pursuing maximum computational speed. This is particularly important for numerical computing since subnormal support enables the gradual underflow mechanism that effectively prevents the sudden loss of precision caused by abrupt underflow, thereby significantly improving numerical stability in long-running iterations, ill-conditioned problems, or scenarios involving the accumulation of very small quantities. Surveys and relevant references can be found in (Goldberg, 1991) and (Higham, 2002).
To show the practical benefit of the support for subnormal numbers, we consider the sequential summation of a geometric series
$\sum_{k=0}^{N-1}s\cdot r^{k}$
where the addends are chosen to straddle the normal-to-subnormal boundary.
As shown in Figure 1, with parameters $N=1000$, $r=0.99$ and $s=2.5\times 10^{-6}$, enabling subnormals allows gradual underflow prevention and disabling them (flush-to-zero) causes early stagnation. This example highlights why faithful IEEE 754 emulation with subnormals is essential for numerical methods involving long accumulations.
Figure 1. Effect of subnormal support in fp16.
@@ 3.2. Fixed-point Arithmetic
Fixed-point is associated with numbers of a fixed number of bits, allocating a predetermined number of bits to the integer part and the fractional part, unlike floating-point representations that use an exponent and significand. The shift to fixed-point arithmetic is driven by several key advantages. First, fixed-point compute units are typically faster and significantly more efficient in terms of hardware resources and power consumption compared to floating-point units. Their smaller logic footprint allows for a higher density of compute units within a given area and power budget. Second, reduced-precision data representation reduces the memory footprint, enabling larger models to fit within available memory while also lowering bandwidth requirements. Fixed-point operations align well with digital signal processors (DSPs) and field-programmable gate arrays (FPGAs) because many lack dedicated floating-point units or optimize for fixed-point arithmetic. Collectively, these benefits enhance data-level parallelism, leading to substantial performance gains (Gupta et al., 2015).
@@
where $I$ is an integer, which may be signed (typically in two’s complement) or unsigned, $f$ is the fixed number of fractional bits, and $n$ is the total number of bits is, with nfn-f bits allocated to the integer part and $f$ bits to the fractional part.
@@
where ibits refers to the bitwidth of integer part, fbits refers to the bitwidth of fractional part, and rmode indicates the rounding mode used in (3.2); the supported rounding modes can be found in Table 3.
The usage example code is given in Appendix A.2.
@@ 3.3. Integer Arithmetic
Integer quantization is a fundamental technique in digital systems to approximate real numbers using a finite set of integers, enabling efficient storage and computation in applications such as machine learning, digital signal processing, and embedded systems. Two primary approaches to quantization exist: symmetric quantization, which balances positive and negative ranges around zero, and asymmetric quantization, which allows unequal ranges for positive and negative values. In principle, integer quantization maps a real number $r\in\mathbb{R}$ to a discrete integer value $x\in\mathbb{Z}$ through scaling and rounding. The process can be adapted for either symmetric or asymmetric quantization depending on the range and scaling strategy. Consider a real number $r$ within a specified range $r_{\text{min}},r_{\text{max}}]$. The goal is to represent $r$ using $n$-bit integers, defining a discrete set of quantization levels. The general quantization process involves scaling, rounding, and clamping, with variations depending on whether symmetric or asymmetric quantization is used.
@@
    For an $n$-bit signed integer in two’s complement, the representable range is $2^{n-1},2^{n-1}-1]$. The quantization range is defined symmetrically as $\omega,\omega]$, where $\omega$ is the maximum absolute value to be represented. The scaling factor $\Delta$ is:
@@
    $\overline{x}=\mathrm{R}(s),\quad\text{ where }\mathrm{R}(\cdot)\text{ denotes general rounding operator.}$
@@
  • Asymmetric Quantization: Asymmetric quantization allows unequal ranges for positive and negative values, typically by defining a range $r_{\text{min}},r_{\text{max}}]$ where $r_{\text{min}}\neq-r_{\text{max}}$. This is useful when the data distribution is skewed (e.g., all positive values in rectified activations like ReLU).
    For an $n$-bit signed integer, the range is still $2^{n-1},2^{n-1}-1]$, but the quantization maps $r_{\text{min}},r_{\text{max}}]$ to this range. The scaling factor $\Delta$ is
@@
    $\overline{x}$ $=\mathrm{R}(s),\quad\text{ where }\mathrm{R}(\cdot)\text{ denotes general rounding operator,}$
@@
    The Chopi framework offers tailored integer quantization, allowing users to specify the bitwidth length (bits) and choose between symmetric or asymmetric quantization (symmetric). Additionally, two channel-specific parameters enable further customization: per_channel determines whether quantization is applied on a per-channel basis, while axis specifies the dimension along which channel-wise quantization occurs when is set to True. A simple code demonstration for integer quantization is given in Appendix A.3.
@@ 3.4. Common Mathematical Functions Support and Array Manipulation Routines
We simulate common mathematical functions and operations (such as built-in functions in NumPy, PyTorch, or JAX) in reduced-precision arithmetic by first rounding the input to low precision, performing operations in the working precision, and then rounding the result back to low precision. This approach contrasts with CPFloat, which applies mathematical operations in working precision to inputs in working precision before rounding the final result to low precision. A usage example is included in Appendix A.4.
Table 4. Functions support and array manipulation routines (part I)44footnotemark: 4
@@
sin Computes sine. Input in radians; output in [1,1][-1,1].
cos Computes cosine. Input in radians; output in [1,1][-1,1].
@@
arcsin Computes arcsin. Input in [1,1][-1,1]; output in $\pi/2,\pi/2]$ radians.
arccos Computes arccos. Input in [1,1][-1,1]; output in $0,\pi]$ radians.
arctan Computes arctan. Output in $\pi/2,\pi/2]$ radians.
@@
tanh Computes hyperbolic tangent. Output in (1,1)(-1,1).
@@
arccosh Computes inverse hyperbolic cosine. Input $\geq 1$; output in $0,\infty)$.
arctanh Computes inverse hyperbolic tangent. Input in (1,1)(-1,1); output real.
@@
log1p Computes $\log(1+x)$. Input >1>-1; enhanced precision for small $x$.
@@
erf Computes error function. Output in (1,1)(-1,1).
@@
Table 5. Functions support and array manipulation routines (part II) [^5] All functions are computed with chopping to enforce reduced-precision format, where applicable.
@@
sign Computes sign. Returns 1-1, 0, or $1$.
clip Clips values to range $a_{\text{min}},a_{\text{max}}]$.